# Hierarchical Sequence Iteration for Heterogeneous Question Answering

### v1

@@ 4 Experiment
HSEQ are evaluated on multiple QA datasets with a focus on both answer quality and efficiency. Metrics include Accuracy, F1, alongside efficiency indicators that reflect the evidence actually inspected: average iteration steps and end-to-end latency (ms).
@@ 4.1 Experiment Setup: Benchmarks and Baselines
To evaluate HSEQ usage from different data modalities, four benchmarks are used for experiments, stressing text-only, table–text hybrid, and KG reasoning: HotpotQA (Yang et al., 2018) (multi-hop reasoning over Wikipedia text), TAT-QA (Zhu et al., 2021a) (table-centric financial QA with accompanying paragraphs and numerical operations), HybridQA (Chen et al., 2020) (Wikipedia tables linked to passages requiring cross-format grounding), and MetaQA (Zhang et al., 2018) over a Wikidata-style KG (Since 1-hop variants are not emphasized due to triviality, during experiment only 2-hop and 3-hop questions are used for experiments).
@@
Three groups are divided for experiments including:
  • LLM-only QA. Multiple LLMs are used to directly answers each question from raw inputs without HSEQ (no unified adapter, no iterative controller), under same prompt instruction.
  • RAG-based methods. Since HSEQ explores different formats of data sources, RAG models specializing in separately Text, Table and Knowledge Graphs have been tested.
@@
  • HSEQ (ours). (i) The best iteration–head pair results and (ii) The median pair results over a grid of open-source models are provided. Three ablations are also included in experiments: (i) LLM-only (no HSEQ); (ii) HSEQ w/o SFT (iteration agent not fine-tuned) and (iii) heuristic-only guidance under fixed template without HSEQ-H.
HSEQ variants.
For the iteration agent (HSEQ-I) and the head agent (HSEQ-H), different LLMs are finetuned and used, listed as:
HSEQ-I: Falcon-H1-3B-Instruct; Qwen3-4B-Instruct-2507; DeepSeek-R1-Distill-Qwen-7B; Falcon3-3B-instruct; Falcon3-7B-instruct; Llama-3.2-3B-Instruct.
HSEQ-H: Falcon3-10B-instruct; Falcon-H1-7B-Instruct; Llama-3.1-8B-Instruct; DeepSeek-R1-Distill-Qwen-7B.
@@
4.2 Experiment Result: How competitive is HSEQ with other baselines?
Table 2 summarizes answer quality across all datasets. HSEQ consistently improves over both LLM-only and strong RAG baselines, while using controlled iteration and exposing explicit provenance. Detailed per-model pairs are reported in Table 3. Efficiency measurements (tokens/latency/steps) are in Table 4.
@@ 4.2 Experiment Result: How competitive is HSEQ with other baselines?
TAT-LLM 74.6 82.9 N/A N/A N/A N/A N/A N/A
TableRAG 47.9 61.9 68.6 N/A N/A N/A N/A N/A N/A
ODYSSEY 51.5 66.0 N/A N/A N/A N/A N/A N/A
TTQA-RS 62.3 70.6 N/A N/A N/A N/A N/A N/A
HippoRAG 65.8 72.4 70.1 74.9 53.2 55.7 N/A N/A N/A N/A
@@
AdaptiveRAG N/A N/A N/A N/A 50.3 52.5 88.2 90.1 84.5 85.7
Our method: HSEQ
HSEQ (best) 66.4 72.1 75.7 83.5 56.3 58.6 95.9 91.1 93.4 88.3
HSEQ (median) 63.9 70.8 73.2 79.6 55.4 57.1 93.2 89.7 90.1 86.6
Our HSEQ achieves strong and consistent gains on multiple benchmarks. On HotpotQA, MetaQA-2hop, and MetaQA-3hop, both the best and median HSEQ configurations surpass all baselines. On TAT-QA, HSEQ’s best run attains the top score overall, while the median run trails slightly behind TAT-LLM (Zhu et al., 2024). On the table-and-text HybridQA, HSEQ attains the best accuracy and the second-best F1 (just behind HippoRAG (Jimenez Gutierrez et al., 2024)); the median configuration remains third among baselines.
@@ 4.3 Yielding between efficiency and accuracy
Table 3 lists results using different HSEQ-I and HSEQ-A. The HybridQA results reveal a clear accuracy–efficiency trade-off across HSEQ agent pairs. The highest accuracy/F1 comes from Qwen3-4B (HSEQ-I) + Falcon-H1-7B (HSEQ-H) (66.2 / 71.4), with the second-best Qwen3-4B + Llama-3.1-8B (65.5 / 71.2). These configurations, however, incur larger iteration depth and latency (about 3.7–4.1 steps; 16.5–21.5 second). On the efficiency end, Llama-3.2-3B + Llama-3.1-8B delivers the lowest steps and latency (2.11; 8.35k ms) with moderate accuracy (55.4 / 57.9), while Falcon3-3B + Falcon-H1-7B attains the second-best efficiency (2.25; 11.7k ms) at similar quality. Taken together, the Pareto frontier spans (i) Qwen-based iterators with larger heads for top accuracy, and (ii) lightweight Llama/Falcon pairs for predictable low latency. Different agent pairs can be chosen regarding whether accuracy or budget dominates.
Table 3: Overall performance of HSEQ agent pairs on Hybrid-QA: Accuracy/F1 and Efficiency.
Accuracy & F1 Efficiency
Iteration Agent (HSEQ-I) Head Agent (HSEQ-H) Avg. Acc Avg. F1 Steps $\downarrow$ Latency (ms) $\downarrow$
Falcon3-3B-instruct Falcon3-10B-instruct 59.3 61.1 2.6 10530.1
Falcon3-3B-instruct Llama-3.1-8B-Instruct 61.2 65.1 2.46 11616.7
Falcon3-3B-instruct Falcon-H1-7B-Instruct 56.1 58.6 2.25 11714.4
@@ 4.4 Efficiency Analysis
To test HSEQ framework’s latency, evidence actually inspected are calculated: iteration steps for HSEQ-I and wall-clock latency are calculated. Results are summarized below. “LLM-only” incurs a single forward pass (1 step) and thus the lowest raw latency, but it lacks provenance and typically underperforms on accuracy in Table 3. HSEQ maintains short, budgeted loops (about 3–5 steps) with substantially lower latency than Graph-centric ToG while preserving multi-hop capability.
@@
HSEQ (ours, best) 4.00 6247.0 3.27 5732.2 4.11 10632.8
HSEQ (ours, median) 4.17 12114.4 3.76 9480.1 4.59 13505.3
4.5 Ablation Studies
Ablation studies are set to evaluate each component of HSEQ framework on representative text (HotpotQA) and table-text (HybridQA) tasks. Following tasks are considered: (a) No SFT (iteration agent not fine-tuned); (b) No guidance (remove $g$); (c) Heuristic-only guidance (no planner) ; and (d) LLM-only (without multi-agent but use HSEQ as part of prompt for data input).
@@ 4.5 Ablation Studies
LLM-only (no HSEQ) 32.9 43.7 21.1 48.3 44.6
The ablation study demonstrates the necessities of all HSEQ’s components, with differing sensitivity across formats. Using heuristic-only guidance yields the smallest degradation from the full system—typically a modest drop in Acc/F1—indicating that a lightweight, template-style prior already guides HSEQ-I effectively when the planner is absent. Removing fine-tuning (w/o SFT) causes a larger decline, but with the use of structured HSEQ data, accuracy remains substantially higher than LLM-only. Without guidance (w/o guidance) influence performance, as in prompt HSEQ-I is only asked to choose necessary evidence from below to answer the question. The results underscore the role of guidance as a portable sufficiency prior. Finally, the LLM-only setting performs worst across all benchmarks, reflecting the difficulty of recovering minimally sufficient evidence without iterative, structure-aware selection. Overall, the results suggest that (i) HSEQ’s unified data structure is the primary source of robustness, (ii) SFT HSEQ-I provides consistent gains, and (iii) guidance—even a simple heuristic ones from template-would increase overall accuracy strongly.
5 Related Work
LLM Finetuning
Large Language Models (LLMs) often adopt finetuning to unlock their capabilities for downstream applications, like medical (Goyal et al., 2024), economic Guo & Yang (2024), or human activity recognition Li et al. (2024). To enhance finetuning efficiency, methods like quantization (Dettmers et al., 2022) parameter efficient fine tuning (Hu et al., 2022; Dettmers et al., 2023; Li & Liang, 2021) can be applied.
Retrieval Augmented Generation
RAG systems help LLMs retrieve extra knowledge according to queries and thereby improving the accuracy of LLM response (Fan et al., 2024), with no necessity to finetune the model. External databases ensure knowledge offered is domain-specific and timely, adding reliability and interpretability (Lewis et al., 2020; Jiang et al., 2023). Accuracy of knowledge retrieval and quality of responses are two key factors for RAG systems evaluation (Yu et al., 2024). Apart from text, table, or html sources (Guo et al., 2024b; Chan et al., 2024; Jin et al., 2025), recent researches have combined graph-structured data into RAG systems(GraphRAG) to improve the efficiency of knowledge interpretability by capturing relationships between entities and utilizing triplets as the primary data source (Edge et al., 2024; Peng et al., 2024; Hu et al., 2024; Mavromatis & Karypis, 2024).
Multi Agent QA system
@@ 6 Conclusion
This paper introduces HSEQ, a compact framework for heterogeneous QA that (i) unifies text, tables, and knowledge graphs into a reversible hierarchical sequence with lightweight structure and provenance; (ii) performs guided, budget-aware iteration that selects small sets of salient segments and predicts sufficiency for early stopping; and (iii) feeds a canonicalized evidence package to a head module for answer synthesis. By replacing single-shot retrieval and unconstrained agentic loops with short, structure-aware selections equipped with an explicit sufficiency signal, HSEQ concentrates computation on evidence actually inspected, delivers predictable latency under token/tool budgets, and preserves auditability through provenance-aware canonicalization.
Across heterogeneous QA benchmarks, HSEQ achieves strong answer quality alongside consistent efficiency, revealing a controllable trade-off between accuracy and cost: larger head with finetuned small iterators achieved both fast and accurate QA. The format-agnostic interface and standardized action schema enable a single learned policy to operate across modalities without per-dataset retrievers, bespoke prompts, or tokenizer changes. Future work will extend HSEQ to multi-turn/streaming settings with dynamic corpora, mitigate hallucination on sufficiency judge under noisy evidence.
7 Ethics Statement.
We affirm adherence to the ICLR Code of Ethics. All experiments use publicly available benchmarks (HybridQA, TatQA, HotpotQA, MetaQA) under their respective licenses; no new human-subject data were collected, and no personally identifiable information (PII) is processed. Our HSEQ construction preserves provenance via identifiers and offsets while avoiding storage of copyrighted text beyond short snippets necessary for QA. As with any LLM-based system, model outputs may reflect societal biases inherited from pretraining corpora; we mitigate this risk by requiring explicit, auditable evidence and by permitting abstention when sufficiency is not met. We release code and configuration solely for research use and discourage deployment in high-stakes settings without domain-specific evaluation and additional safeguards (fairness, privacy, and safety audits).
8 Reproducibility Statement.
We provide an anonymous GitHub link (https://anonymous.4open.science/r/HSEQ-anonymous-0DAC) with code and scripts to (i) construct HSEQ from raw corpora, (ii) fine-tune the iteration policy with LoRA, and (iii) run guided inference and evaluation. Implement details are shown in App. A.3, containing models used (App. A.3.1), prompts (App. A.3.2A.3.3), LoRA adaption parameters (App. A.3.4) and reproducibility notes (App. A.3.6). Theorems include complete assumptions and proofs (App. A.1). Apart from the code, detailed examples of agents interactions (example questions, LLM outputs, data retreived each steps, etc.) are provided in App. A.5 and as a jsonl file in our anonymous repository.
Acknowledgments
This research is partially support by Technology Innovation Institure Abu Dhabi, UAE. Also, this research includes computations using the computational cluster supported by SHARON AI at Sydney, New South Wales.
@@ Appendix A Appendix
A.1 Theoretical Properties of HSEQ
@@ A.1.3 Windowed Iteration: Coverage and Complexity
Let $C(W)>0$ be the (deterministic) per-step context cost determined by window size $W$. Under admissibility, the total selection cost is bounded by
$\mathrm{Cost}_{\mathrm{select}}\ \leq\ C(W)\cdot\min\{T_{\max},\lceil L/k\rceil\},$
independent of $|S_{h}|$. If $L$ is a nonnegative integer random variable with $\mathbb{E}[L]=\bar{L}<\infty$, then
$\mathbb{E}\big[\mathrm{Cost}_{\mathrm{select}}\big]\ \leq\ C(W)\cdot\mathbb{E}\!\left[\min\{T_{\max},\lceil L/k\rceil\}\right]\ \leq\ C(W)\cdot\min\{T_{\max},\,\bar{L}/k+1\}.$
@@ A.3 Implementation Details
A.3.1 Agent models used for HSEQ
@@ A.3.5 Canonicalization and Sufficiency
At termination, a modality-agnostic canonicalizer $\kappa$ converts the selected set $M_{\tau}$ into a compact, auditable structure
$\kappa(M_{\tau})\;=\;\big\{(\texttt{id},\texttt{level},\texttt{uri},\texttt{offsets},\texttt{source\_type},\texttt{snippet}\,;\,\texttt{meta})\big\}_{s\in M_{\tau}},$
with the following contract: (i) id: globally unique, deterministically derived (e.g., $\mathrm{sha1}(\texttt{uri},\texttt{offsets})$); (ii) uri: source identifier with version (e.g., document path or graph name); (iii) offsets: zero-based half-open character indices $[a,b)$ into the original source; for tables, $[i,j]$ denotes row/column coordinates; for KGs, offsets=(1,1)=(-1,-1); (iv) snippet: a human-readable content aligned to sentence/field boundaries when possible; (v) meta: integrity and alignment helpers (schema, time, source_version, sha1). Duplicates are removed by $(\texttt{uri},\texttt{offsets})$ and the package is deterministically ordered by uri then offsets. Typed views are derived on demand: text $\Rightarrow$ spans with section/paragraph ids; table $\Rightarrow$ row_id, col_ids, schema, cell_coords; KG $\Rightarrow$ $(h,r,t)$ plus optional validity time.
@@
The sufficiency head outputs $s_{t}\in\{0,1\}$ at each step. Training targets follow a coverage-based heuristic: $s_{t}^{\star}=1$ if and only if the current $M_{t}$ satisfies task-specific adequacy (e.g., contains at least one gold-positive segment; achieves full slot coverage for table QA; or yields a unique answer span/number under a fixed head). For weak supervision, per-step weights downweight low-confidence positives (App. A.2). Inference uses a calibrated threshold $\tau$ on the model’s sufficiency score $\hat{p}_{t}$ and enforces a minimum step count $T_{\min}$:
$\text{stop at }\tau\;=\;\min\{t\geq T_{\min}:\,\hat{p}_{t}\geq\tau\}\quad\text{or when budget }B\text{ is exhausted.}$
Optionally, a lightweight contradiction checker triggers a one-shot refinement loop of at most $\Delta$ additional steps with tightened guidance $g^{\prime}$ and reduced budget $B^{\prime}$. Thresholds $(\tau,T_{\min})$ are selected on the development split and may be calibrated via temperature scaling.

### v3

@@ 5 Experiment
RELOOP are evaluated on multiple QA datasets with a focus on both answer quality and efficiency. Metrics include Accuracy, F1, alongside efficiency indicators that reflect the evidence actually inspected: average iteration steps and end-to-end latency (ms).
@@ 5.1 Experiment Setup: Benchmarks and Baselines
To evaluate RELOOP usage from different data modalities, four benchmarks are used for experiments, stressing text-only, table–text hybrid, and KG reasoning: HotpotQA (Yang et al., 2018) (multi-hop reasoning over Wikipedia text), TAT-QA (Zhu et al., 2021a) (table-centric financial QA with accompanying paragraphs and numerical operations), HybridQA (Chen et al., 2020) (Wikipedia tables linked to passages requiring cross-format grounding), and MetaQA (Zhang et al., 2018) over a Wikidata-style KG (Since 1-hop variants are not emphasized due to triviality, during experiment only 2-hop and 3-hop questions are used for experiments).
@@
Three groups are divided as:
  • LLM-only QA. Multiple LLMs are used to directly answers each question from raw knowledge source inputs, under same prompt instruction.
  • RAG-based methods. Since RELOOP explores different formats of data sources, RAG models specializing in separately Text, Table and Knowledge Graphs have been tested.
@@
  • RELOOP (ours). (i) The best iteration–head pair results and (ii) The median pair results over a grid of open-source models are provided. Three ablations are also included in experiments: (i) LLM-only ; (ii) RELOOP w/o SFT (RELOOP-I not fine-tuned) (iii) without guidance from RELOOP-Hand (iii) with only guidance from template.
RELOOP variants.
For the iteration agent (RELOOP-I) and the head agent (RELOOP-H), different LLMs are finetuned and used, listed as:
RELOOP-I: Falcon-H1-3B-Instruct; Qwen3-4B-Instruct-2507; DeepSeek-R1-Distill-Qwen-7B; Falcon3-3B-instruct; Falcon3-7B-instruct; Llama-3.2-3B-Instruct.
RELOOP-H: Falcon3-10B-instruct; Falcon-H1-7B-Instruct; Llama-3.1-8B-Instruct; DeepSeek-R1-Distill-Qwen-7B.
@@
TAT-LLM (Table task) N/A N/A 73.1 81.0 N/A N/A N/A N/A N/A N/A
TableRAG (Table task) 47.9 61.9 68.6 N/A N/A N/A N/A N/A N/A
ODYSSEY (Text task) 51.5 66.0 N/A N/A N/A N/A N/A N/A N/A N/A
TTQA-RS (Text task) 62.3 70.6 N/A N/A N/A N/A N/A N/A N/A N/A
HippoRAG (Table + text) 65.8 72.4 70.1 74.9 53.2 55.7 N/A N/A N/A N/A
@@
AdaptiveRAG (Graph task) N/A N/A N/A N/A 50.3 52.5 88.2 90.1 84.5 85.7
Our method: RELOOP
RELOOP (best) 66.4 72.1 75.7 83.5 56.3 58.6 95.9 91.1 93.4 88.3
RELOOP (median) 63.9 70.8 73.2 79.6 55.4 57.1 93.2 89.7 90.1 86.6
5.2 Experiment Result: How competitive is RELOOP with other baselines?
Table 2 summarizes answer quality across all datasets. RELOOP consistently improves over both LLM-only and strong RAG baselines, while using controlled iteration and exposing explicit provenance. Detailed per-model pairs are reported in Table 8. Efficiency measurements (tokens/latency/steps) are in Table 3.
Our RELOOP achieves strong and consistent gains on multiple benchmarks. On HotpotQA, MetaQA-2hop, and MetaQA-3hop, both the best and median RELOOP configurations surpass all baselines. On TAT-QA, RELOOP’s best run attains the top score overall, while the median run trails slightly behind TAT-LLM (Zhu et al., 2024). On the table-and-text HybridQA, RELOOP attains the best accuracy and the second-best F1 (just behind HippoRAG (Jimenez Gutierrez et al., 2024)); the median configuration remains third in baselines.
@@ 5.3 Yielding between efficiency and accuracy
Table 8 in App. A.5 lists results using different RELOOP-I and HSEQ-A. The HybridQA results reveal a clear accuracy–efficiency trade-off across RELOOP agent pairs. The highest accuracy/F1 comes from Qwen3-4B (RELOOP-I) + Falcon-H1-7B (RELOOP-H) (66.2 / 71.4), with the second-best Qwen3-4B + Llama-3.1-8B (65.5 / 71.2). These configurations, however, incur larger iteration depth and latency (about 3.7–4.1 steps; 16.5–21.5 second). On the efficiency end, Llama-3.2-3B + Llama-3.1-8B delivers the lowest steps and latency (2.11; 8.35k ms) with moderate accuracy (55.4 / 57.9), while Falcon3-3B + Falcon-H1-7B attains the second-best efficiency (2.25; 11.7k ms) at similar quality. Taken together, the Pareto frontier spans (i) Qwen-based iterators with larger heads for top accuracy, and (ii) lightweight Llama/Falcon pairs for predictable low latency. Different agent pairs can be chosen regarding whether accuracy or budget dominates.
@@ 5.4 Efficiency Analysis
To test RELOOP framework’s latency, evidence actually inspected are calculated: iteration steps for RELOOP-I and wall-clock latency are calculated. Results are summarized below. “LLM-only” incurs a single forward pass (1 step) and thus the lowest raw latency, but this comes at the cost of weaker multi-hop accuracy and no explicit provenance in Table 8. In contrast, graph-centric ToG performs many expansion steps (11–17 on average), which substantially increases latency (e.g., over 2222k ms on HotpotQA and 2424k ms on MetaQA-3hop), even though it is designed for multi-hop reasoning.
@@
RELOOP (ours, best) 4.00 6247.0 3.27 5732.2 4.11 10632.8
RELOOP (ours, median) 4.17 12114.4 3.76 9480.1 4.59 13505.3
@@
LLM-only 32.9 43.7 21.1 48.3 44.6
RELOOP occupies a middle ground in this trade-off. Both the best and median RELOOP variants maintain short, budgeted loops of roughly 3355 steps across datasets, yet reduce latency by more than half relative to ToG on all three benchmarks. This indicates that guided, windowed iteration over RELOOP can retain multi-hop capability while avoiding the long expansion chains and repeated graph traversals of ToG. Compared with LLM-only, RELOOP pays a moderate overhead in latency but gains structured evidence and substantially higher accuracy on multi-step questions. RELOOP provides a more balanced operating point with bounded steps and competitive performance.
5.5 Ablation Studies
Ablation studies are set to evaluate each component of RELOOP framework on representative text (HotpotQA) and table-text (HybridQA) tasks. Following tasks are considered: (a) No SFT (RELOOP-I not fine-tuned); (b) No guidance (remove $g$); (c) Heuristic-only guidance (no planner) ; and (d) LLM-only (without multi-agent but use HSEQ as data input).
The ablation study demonstrates the necessities of all RELOOP’s components, with differing sensitivity across formats. Using heuristic-only guidance yields the smallest degradation from the full system—typically a modest drop in Acc/F1—indicating that a lightweight, template-style prior already guides RELOOP-I effectively when the planner is absent. Removing fine-tuning (w/o SFT) causes a larger decline, but with the use of structured HSEQ data, accuracy remains substantially higher than LLM-only. Without guidance (w/o guidance) influence performance, as in prompt RELOOP-I is only asked to choose necessary evidence from below to answer the question. The results underscore the role of guidance as a portable sufficiency prior. Finally, the LLM-only setting performs worst across all benchmarks, reflecting the difficulty of recovering minimally sufficient evidence without iterative, structure-aware selection. Overall, the results suggest that (i) RELOOP’s unified data structure is the primary source of robustness, (ii) SFT RELOOP-I provides consistent gains, and (iii) guidance—even a simple heuristic ones from template-would increase overall accuracy strongly.
@@ 6 Conclusion
This paper introduces RELOOP, a compact framework for heterogeneous QA that (i) unifies text, tables, and knowledge graphs into a reversible hierarchical HSEQ with lightweight structure and provenance; (ii) performs guided, budget-aware iteration that selects small sets of salient segments and predicts sufficiency for early stopping; and (iii) feeds a canonicalized evidence package to a head module for answer synthesis. By replacing single-shot retrieval and unconstrained agentic loops with short, structure-aware selections equipped with an explicit sufficiency signal, RELOOP concentrates computation on evidence actually inspected, delivers predictable latency under token/tool budgets, and preserves auditability through provenance-aware canonicalization.
Across heterogeneous QA benchmarks, RELOOP achieves strong answer quality alongside consistent efficiency, revealing a controllable trade-off between accuracy and cost: larger head with finetuned small iterators achieved both fast and accurate QA. The unified sequence parsing and standardized action schema enable a single learned policy to operate across modalities reliably.
Limitations
Although through experiments are done to prove RELOOP’s universality under different modalities of data, more improvement can be added on the framework. Future work will extend RELOOP to multi-turn/streaming settings with dynamic corpora, mitigate hallucination on sufficiency judge under noisy evidence.
@@ Appendix A Appendix
A.1 Theoretical Properties of RELOOP
@@ A.1.3 Windowed Iteration: Coverage and Complexity
Let $C(W)>0$ be the (deterministic) per-step context cost determined by window size $W$. Under admissibility, the total selection cost is bounded by $\mathrm{Cost}_{\mathrm{select}}\ \leq\ C(W)\cdot\min\{T_{\max},\lceil L/k\rceil\},$ independent of $|S_{h}|$. If $L$ is a nonnegative integer random variable with $\mathbb{E}[L]=\bar{L}<\infty$, then $\mathbb{E}\big[\mathrm{Cost}_{\mathrm{select}}\big]\ \leq\ C(W)\cdot\mathbb{E}\!\left[\min\{T_{\max},\lceil L/k\rceil\}\right]\ \leq\ C(W)\cdot\min\{T_{\max},\,\bar{L}/k+1\}.$
@@ A.3 Implementation Details
A.3.1 Agent models used for RELOOP
@@ A.3.4 LoRA Adaptation and Optimization
Table 6 shows the SFT hyperparameters for each model-size group.
@@ A.3.5 Canonicalization and Sufficiency
At termination, a modality-agnostic canonicalizer $\kappa$ converts $M_{\tau}$ into a compact record set: $\kappa(M_{\tau})=\{\,r_{s}\,\}_{s\in M_{\tau}},\qquad r_{s}=(\texttt{id},\texttt{lvl},\texttt{uri},\texttt{off},\texttt{typ},\texttt{snip},\texttt{meta}).$
where lvl=level, off=offsets, typ=source_type, and snip=snippet. with the following contract: (i) id: globally unique, deterministically derived (e.g., $\mathrm{sha1}(\texttt{uri},\texttt{offsets})$); (ii) uri: source identifier with version (e.g., document path or graph name); (iii) offsets: zero-based half-open character indices $[a,b)$ into the original source; for tables, $[i,j]$ denotes row/column coordinates; for KGs, offsets=(1,1)=(-1,-1); (iv) snippet: a human-readable content aligned to sentence/field boundaries when possible; (v) meta: integrity and alignment helpers (schema, time, source_version, sha1). Duplicates are removed by $(\texttt{uri},\texttt{offsets})$ and the package is deterministically ordered by uri then offsets. Typed views are derived on demand: text $\Rightarrow$ spans with section/paragraph ids; table $\Rightarrow$ row_id, col_ids, schema, cell_coords; KG $\Rightarrow$ $(h,r,t)$ plus optional validity time.
@@
The sufficiency head outputs $s_{t}\in\{0,1\}$ at each step. Training targets follow a coverage-based heuristic: $s_{t}^{\star}=1$ if and only if the current $M_{t}$ satisfies task-specific adequacy (e.g., contains at least one gold-positive segment; achieves full slot coverage for table QA; or yields a unique answer span/number under a fixed head). For weak supervision, per-step weights downweight low-confidence positives (App. A.2). Inference uses a calibrated threshold $\tau$ on the model’s sufficiency score $\hat{p}_{t}$ and enforces a minimum step count $T_{\min}$: $\text{stop at }\tau\;=\;\min\{t\geq T_{\min}:\,\hat{p}_{t}\geq\tau\}\quad\text{or when budget }B\text{ is exhausted.}$ Optionally, a lightweight contradiction checker triggers a one-shot refinement loop of at most $\Delta$ additional steps with tightened guidance $g^{\prime}$ and reduced budget $B^{\prime}$. Thresholds $(\tau,T_{\min})$ are selected on the development split and may be calibrated via temperature scaling.